What is the future value of $600 in 21 years assuming an interest rate of 11 percent compounded semiannually?

Multiple Choice

  • $842.98

  • $737.89

  • $5,685.32

  • $5,369.50

  • $5,401.05
Answer #1

$5,685.32

Future Value of investment = Investment Amount * Future Value of 1
= $           600 * 9.475525
= $ 5,685.32
Working:
Future Value of 1 = (1+i)^n Where,
= (1+0.055)^42 i = 11%/2 = 0.055
= 9.4755255 n = 21*2 = 42